OpenStudy (snuggielad):

Surprise! I'm actually asking a course related question! Simple problem, maybe it's because I'm tired right now but I don't even know where to start! Help is appreciated!

OpenStudy (snuggielad):

Identify the set of values for X for which Y will be a real number. \[y=\frac{ 2 }{ x + 7 }\]

OpenStudy (btaylor):

We want y to be a real number, so 2/(x+7) has to be rational. What value(s) of x are impossible?

OpenStudy (snuggielad):

-5? But isn't this asking for a set?

OpenStudy (btaylor):

No, -5 would still work. You'd get 2/2, which is definitely real. If you have division, the denominator of the fraction can't be zero (because it will be undefined).

OpenStudy (snuggielad):

Identify the "set of values" for X for which Y will be a real number. Set, meaning multiple. Values, being plural also meaning multiple. It's not asking for "a value of x" for which y will be a real number. How would I come up with said set?

imqwerty (imqwerty):

it is not necessary that for \(\Large\frac{2}{x+7}\) to be real it must have to be rational. well any real number is a number that can be represented on the number like. rational numbers like 2,7,5.. are all real irrational numbers like \(\sqrt{2}\),\(\sqrt{3}\).. are also real (because they can be represented on number line). this might help-> \(\Large\frac{any~real}{any~real}=real\)

imqwerty (imqwerty):

but remember that indeterminate form like \(\Large \frac{some~number}{0}\) is not considered to be real

OpenStudy (snuggielad):

Ah, gotcha.

OpenStudy (snuggielad):

So, for example, same directions but different question: \[y=\sqrt{x+11}\]

OpenStudy (snuggielad):

I would assume any positive?

OpenStudy (snuggielad):

Bad grammar I apologize. I would assume, Any positive?*

imqwerty (imqwerty):

yes in this question the thing inside the square root cannot be negative so \((x+11)\) must be positive

OpenStudy (snuggielad):

Thanks for your help! I think I was just confused as to the wording of the question. God Bless! ~SnuggieLad

imqwerty (imqwerty):

Lol np :D

OpenStudy (snuggielad):

Now, where did I put my pencil? ;) I hate it when that happens.

imqwerty (imqwerty):

yeah happens ;) normally find it kept inside some book eh xd i got an exam in 20min gtg :D